### v1

# A Simple Derived Categorical Generalization of Ulrich Bundles

Abstract.
We define special objects, Ulrich objects, on a derived category of polarized smooth projective variety $(X,\mathscr{O}(1))$ as a generalization of Ulrich bundles to the derived category. These are defined by the cohomological conditions that are the same form as a cohomological criterion determining Ulrichness for sheaves.
This paper gives a characterization of the Ulrich object similar to the one in [ES03]. As an application, we have provided a new approach to the Eisenbud-Schreyer question by using the notions of the generator of the derived category. We also have given an example of Ulrich objects that are not sheaf by the Yoneda extension.
Key words and phrases:
Derived category, Ulrich bundles, Generators of triangulated categories, Algebraic curves, Bridgeland stability condition
2020 Mathematics Subject Classification:
14F08 (primary), 14J60, 14D20 (secondary).
0. Introduction
Let $X$ be a smooth projective variety over $k=\overline{k}$ and $D^{b}(X)$ be the bounded derived category of coherent sheaves on $X$. A derived category, whose objects are bounded complexes of coherent sheaves, is a triangulated category constructed from an abelian category. The investigation of $D^{b}(X)$ as an invariant of variety began with the work of Bondal and Olrov [BO01], where they showed that if $X$ has (anti-)ample canonical line bundle, then the equivalence $D^{b}(X)\cong D^{b}(Y)$ implies $X\cong Y$.
Since then, the derived category has been extensively studied in relation to the minimal model program [Kaw02] and Bridgeland stability conditions [Bri09], which is a generalized notion of sheaf stability.
Meanwhile, recent researches has shown that the derived categories are also valuable to investigate the properties of sheaves. For instance, derived categorical methods are employed in the works of Ohno [Ohn20, Ohn22] and Fukuoka, Hara, and Ishikawa [FHI22, FHI23] to classify certain classes of special bundles.
This paper proposes that the derived categorical framework is likewise valuable for the study of Ulrich bundles, particularly regarding the existence problem, which has attracted the attention of algebraic geometers in recent years.
0.1. Ulrich Bundles
Let $X\subset\mathbb{P}^{N}$ be a smooth projective variety embedded into the projective space $\mathbb{P}^{N}$. Let $\mathscr{O}_{X}(1)\coloneqq\mathscr{O}_{\mathbb{P}^{N}}(1)\otimes\mathscr{O}_{X}$. A sheaf $\mathcal{E}$ on $X$ is called an Ulrich sheaf if it is initialized, is arithmetically Cohen-Macaulay, and satisfies vanishing conditions of intermediate cohomologies with respect to $\mathscr{O}_{X}(1)$ (see 1.3).
In [Ulr84], Ulrich introduced the notion of Ulrich bundles in the context of commutative algebra. Motivated by Beauville’s results [Bea00], Eisenbud and Schreyer [ES03] later brought this notion to algebraic geometry. One of the primary issues regarding Ulrich bundles is their existence. Eisenbud and Schreyer posed the following question, which is now widely believed to have a positive answer and is known as Eisenbud–Schreyer conjecture.
Question A ([ES03]).
Is every smooth projective variety $X\subset\mathbb{P}^{N}$ the support of an Ulrich sheaf?
The existence problem has been studied for some varieties through ad hoc methods. For example, the Ulrich bundles on the projective spaces are given by (direct sums of) structure sheaves, and these on the quadric hypersurface $Q^{n}\subset\mathbb{P}^{n+1}$ by Spinor bundle(s). For more general hypersurfaces, their existences are studied via matrix factorizations. Additionally, the proof of existence often relies on the choice of a very ample divisor on $X$.
0.2. Ulrich Objects
In this paper, we introduce the notion of Ulrich objects as a generalization of Ulrich bundles in terms of the derived category. These objects are defined by cohomological conditions analogous to the cohomological criterion of Ulrichness for sheaves (see 2.1). Firstly, we begin with establishing a characterization of Ulrich objects that is partially parallel to the characterization of Ulrich bundles (1.5).
Let $X$ be a smooth projective variety with an embedding $X\hookrightarrow\mathbb{P}^{N}$ by a linear system $|\mathscr{O}_{X}(1)|$ and $\pi:X\to\mathbb{P}^{\dim X}$ a finite linear projection (1.4).
Theorem B (see 2.2).
With the setting above, the following conditions are equivalent:
  1. (1)
    $E\in D^{b}(X)$ is an Ulrich object,
  2. (2)
    $\pi_{\ast}E\in\langle\mathscr{O}\rangle$,
  3. (3)
    each cohomology sheaf $\mathcal{H}^{i}(E)$ is an Ulrich bundle if it is not zero.
The last condition in the Theorem seems to mean that one does not get something new by this generalization. However, the last condition in B plays an important role in the study of the existence of Ulrich objects and bundles. Namely, the existence of the Ulrich object is equivalent to the existence of the Ulrich bundle.
Also, we present an example of an Ulrich object that is not an Ulrich bundle: a Yoneda type Ulrich object.
0.3. Application to the Eisenbud-Schreyer’s Question
As stated above, to establish the existence of Ulrich bundles on $(X,\mathscr{O}(1))$, it suffices to show the existence of Ulrich objects. In terms of the derived category, the existence of the Ulrich object is equivalent to the following claim:
$\mathcal{G}^{\prime}\coloneqq\mathscr{O}_{X}(1)\oplus\cdots\oplus\mathscr{O}_{X}(n)\text{ is not a generator}.$
In Section 4.1, we will review the definitions of various types of generators in triangulated categories, as well as the relationship among them.
We present a new general proof of the existence of Ulrich bundles on an algebraic curve with an arbitrary embedding, based on the framework of the derived category. Note that the existence of Ulrich line bundle on a curve has already established in [ES09]. To approach the existence problem, our proof is divided into two parts. The first part is the following proposition:
Proposition C (see 4.9).
Let $\mathcal{A}$ be a hereditary abelian category and $\mathcal{D}(\mathcal{A})$ the derived category of $\mathcal{A}$ with arbitrary direct sums. If $G\in\mathcal{D}^{\mathrm{cpt}}(\mathcal{A})$ is a weak generator of $\mathcal{D}^{\mathrm{cpt}}(\mathcal{A})$, then $G$ is also a generator of $\mathcal{D}(\mathcal{A})$.
The second step is showing that $\mathcal{G}^{\prime}$ cannot classically generate $D^{b}(C)$. By combining B and C, we arrive at the following corollary:
Theorem D (see 4.10).
For any smooth projective curve with an embedding $(C,\mathscr{O}_{C}(1))$, there exists an Ulrich object.
Related Work and Further Questions.
The existence problem has been widely studied and there are too many references to list them out here, but we mention some of them:
As stated above, the curve case treated in [ES09]. The results for hypersurfaces and, more generally, complete intersections are studied in [HUB91, BHU87]. Also, Grassmannians are discussed in [CMR15]. [FP23] discusses the property of Ulrich bundles via the Bridgeland stability conditions.
We believe this method can be extended to more general cases, potentially providing a comprehensive answer to the A.
Outline of this paper.
Section 1 reviews the preliminary results on Ulrich bundles and some basic notions of the bounded derived category of coherent sheaves for clarifying the definition of the Ulrich object. The Ulrich object will be introduced in Section 2. Also, Section 2 contains the proof of B.
The following Section 3 discusses the example of Ulrich objects. In particular, we introduce a special type of Ulrich object, called Yoneda type Ulrich object. The application of Ulrich object to A is given in Section 4. Especially, the proof of D is in Section 4.2. Finally, some remarks about the relationship between the Ulrich objects and (divisorial) Bridgeland stability conditions are in Section 5.
Notation and Conventions.
The convention and notations used in this paper are listed below:
  • all varieties are smooth and projective defined over $k=\overline{k}$,
  • the dimension of the variety $X$ is written by $n$ unless otherwise stated,
  • $\mathcal{D}$ represents a triangulated category with arbitrary direct sums,
  • $\mathcal{D}^{\mathrm{cpt}}$ is a full subcategory of $\mathcal{D}$ that consists of compact objects,
  • for an abelian category $\mathcal{A}$, $D^{b}(\mathcal{A})$ is the bounded derived category of $\mathcal{A}$,
  • the bounded derived category $D^{b}(\operatorname{\mathbf{Coh}}(X))$ is denoted by $D^{b}(X)$ for short,
  • $D(\operatorname{\mathbf{Qcoh}}(X))$ is the unbounded derived category of quasi-coherent sheaves on the scheme $X$,
  • $\operatorname{Perf}(X)$ is a triangulated category consisting of perfect complexes on $X$ (when $X$ is smooth, $D(\operatorname{\mathbf{Qcoh}}(X))^{\textrm{cpt}}=\operatorname{Perf}(X)=D^{b}(X)$), and
  • all functors are derived unless otherwise stated. However, $\mathbf{L}$ and $\mathbf{R}$ are occasionally used to emphasize the functor is actually derived.
Acknowledgement.
The aurhor is grateful to his advisor Professor Yasunari Nagai for his continuous encouragement and fruitful discussions. He would also like to thank Professors Wahei Hara and Yuki Mizuno for useful discussions. The author was supported by JST SPRING, Grant Number JPMJSP2128.
1. Towards the Definition of Ulrich Object
This Section 1 contains the preliminary results of Ulrich bundles and derived categories of coherent sheaves for non-experts of each topic. Thus, there is no problem for experts to skip Section 1.
1.1. Preliminaries on Ulrich Bundles
Most of the basic results of Ulrich bundles are cited from the book about Ulrich bundles [CMRPL21]. Also, [Bea18] is an article for an introduction to Ulrich bundles.
Definition 1.1.
A sheaf $\mathcal{E}$ on $X$ is called arithmetically Cohen-Macaulay (aCM for short) if locally Cohen-Macaulay (i.e. $\operatorname{depth}\mathcal{E}_{x}=\dim\mathscr{O}_{X,x}$ for any $x\in X$) and $H^{i}(X,\mathcal{E}(t))=0$ for any $t\in\mathbb{Z}$ and $i=1,\dots,n-1$.
Definition 1.2.
A sheaf $\mathcal{E}$ on $X$ is called initialized if $H^{0}(X,\mathcal{E})\neq 0$ and $H^{0}(X,\mathcal{E}(t))=0$ for any $t<0$.
Definition 1.3.
A sheaf $\mathcal{E}$ on $X$ is called Ulrich if it is initialized aCM and $h^{0}(X,\mathcal{E})=\deg X\operatorname{rk}\mathcal{E}$.
Remark.
If $\mathcal{E}$ is initialized and aCM, the inequality holds;
$h^{0}(\mathcal{E})\leq\deg(X)\operatorname{rk}\mathcal{E}.$
Remark.
Any aCM bundle is a vector bundle. Indeed, by the Auslander-Buchsbaum formula,
$\operatorname{pd}\mathcal{E}_{x}=\operatorname{depth}\mathscr{O}_{X,x}-\operatorname{depth}\mathcal{E}_{x}=0.$
Thus, $\mathcal{E}_{x}$ is $\mathscr{O}_{X,x}$- free module for any $x\in X$, noting that the smoothness of the variety is used here. Thus, Ulrich sheaves are sometimes referred to as Ulrich bundles.
Definition 1.4.
For an polarized variety $(X,\mathscr{O}(1))$, the morphism $\pi:X\to\mathbb{P}^{n}$ as following diagram:
${X}$${\mathbb{P}^{N}}$${\mathbb{P}^{n},}$$\scriptstyle{\Phi_{|\mathscr{O}(1)|}}$$\scriptstyle{\pi}$$\scriptstyle{p}$
where the $p$ is the composition of linear projection from the points in $\mathbb{P}^{N}\setminus X$. The composition $\pi\coloneqq p\circ\Phi_{|\mathscr{O}(1)|}$ that is defined above is called finite linear projection.
Remark.
The finite linear projection is a finite morphism. Indeed, as the linear projections are quasi-finite and proper, this is a finite morphism.
The motivation for considering the Ulrich bundles goes back to the paper [Bea00]. Let $X\subset\mathbb{P}^{n+1}$ be a hypersurface of degree $d$, defined by the equation $F=0$. Then, Beauville showed that $X$ behaves like a determinantal variety in the meaning of $F^{r}=\det M$ for some $r$ and $M$ if and only if there exists a rank $r$ vector bundle $\mathcal{E}$ on $X$ that admits a linear resolution
$0\to\mathscr{O}_{\mathbb{P}^{n+1}}(-1)^{\oplus rd}\stackrel{{\scriptstyle M}}{{\longrightarrow}}\mathscr{O}_{\mathbb{P}^{n+1}}^{\oplus rd}\to\mathcal{E}\to 0$
Generalizing this result, Eisenbud and Schreyer have shown that the following result:
Theorem 1.5 ([ES03, Proposition 2.1.]).
Let $\mathcal{E}$ be an initialized vector bundle on $n$-dimensional variety $X$ and an embedding $\Phi_{|\mathscr{O}_{X}(1)|}:X\hookrightarrow\mathbb{P}^{N}$. The following statements are equivalent:
  1. (1)
    $\mathcal{E}$ is an Ulrich bundle,
  2. (2)
    There is a linear resolution
    $0\to\mathscr{O}_{\mathbb{P}^{N}}(-N+n)^{\oplus a_{N-n}}\to\cdots\to\mathscr{O}_{\mathbb{P}^{N}}(-1)^{\oplus a_{1}}\to\mathscr{O}_{\mathbb{P}^{N}}^{\oplus a_{0}}\to\mathcal{E}\to 0$
  3. (3)
    $H^{i}(\mathcal{E}(-j))=0$ for any $i\geq 0$ and $j=1,\dots,n$,
  4. (4)
    For some finite linear projection $\pi:X\to\mathbb{P}^{n}$, $\pi_{\ast}\mathcal{E}\cong\mathscr{O}^{\oplus m}_{\mathbb{P}^{n}}$ for some $m$.
The rest of this subsection is devoted to reviewing the stability of Ulrich bundles.
Theorem 1.6 ([CHGS12, Theorem 2.9.]).
Let $(X,\mathscr{O}(1))$ be a $n$-dimensional variety and $E$ be an Ulrich bundle on $X$. Then, $E$ is semistable in the sense of Gieseker (In particular, $\mu$-semistable) with respect to $\mathscr{O}(1)$.
Proposition 1.7.
Let
$0=\mathcal{E}_{0}\subset\cdots\subset\mathcal{E}_{n-1}\subset\mathcal{E}_{n}=\mathcal{E}$
be a Jordan-Hölder filtration of Ulrich bundle $\mathcal{E}$. Then, each sub-sheaf $\mathcal{E}_{i}$ is also an Ulrich bundle.
1.2. Concise reviews on Derived Categories
In Section 1.2, we shortly give reviews on some definitions and notations of derived categories of coherent sheaves for non-experts of derived categories. The derived category of an abelian category has the structure of a triangulated category. Some of the following results can be applied to triangulated categories. However, we will only mention the result for derived categories for simplicity.
Let $X$ be a smooth projective variety of dimension $n$ over $\mathbb{C}$, and $D^{b}(X)\coloneqq D^{b}(\operatorname{\mathbf{Coh}}(X))$ be the bounded derived category of coherent sheaves. The object $E\in D^{b}(X)$ is, as an entity, the complex of quasi-coherent sheaves on $X$ with coherent cohomologies like:
$E:\big{(}\cdots\to E_{-1}\to E_{0}\to E_{1}\to E_{2}\to\cdots\big{)}.$
The $i$-th cohomology sheaf of $E$ as a complex is written as $\mathcal{H}^{i}(E)\in\operatorname{\mathbf{Coh}}(X)$ and $H^{i}(E)\coloneqq\mathcal{H}^{i}(\mathbf{R}\Gamma(E))\in\mathbf{Vec}$, here $\mathbf{R}\Gamma:D^{b}(X)\to D^{b}(\mathbf{Vec})$ is the derived functor of global section functor.
Definition 1.8.
Let $\mathcal{C}\subset\mathcal{D}$ be a full subcategory or a set of objects.
  1. (1)
    The thick closure of $\mathcal{C}$ in $\mathcal{D}$ is the smallest full triangulated subcategory containing $\mathcal{C}$ which is closed under taking direct summands. Denote this by $\operatorname{thick}\mathcal{C}$.
  2. (2)
    The full subcategory $\mathcal{C}$ is called thick if $\mathcal{C}=\operatorname{thick}\mathcal{C}$ holds. Denote the smallest thick subcategory containing $\mathcal{C}$ by $\langle\mathcal{C}\rangle$.
Definition 1.9.
Let $\mathcal{D}$ be a triangulated category. A sequence of full subcategories $\{\mathcal{C}_{1},\ldots,\mathcal{C}_{l}\}$ forms a semi-orthogonal decomposition of $\mathcal{D}$ if the following two conditions hold:
  1. (1)
    $\mathop{\mathrm{Hom}}\nolimits(\mathcal{C}_{m},\mathcal{C}_{n})=0$ for $m>n$,
  2. (2)
    for any object $D\in\mathcal{D}$, there is a sequence
    $0=D_{l}\rightarrow D_{l-1}\rightarrow\cdots\rightarrow D_{1}\rightarrow D_{0}=D,$
    such that $\operatorname{Cone}(D_{i}\rightarrow D_{i-1})\in\mathcal{C}_{i}$.
Denote the semi-orthogonal decomposition by $\mathcal{D}=\langle\mathcal{C}_{1},\ldots,\mathcal{C}_{l}\rangle$ (Indeed, in this case, the thick closure of $\mathcal{C}_{1},\ldots,\mathcal{C}_{l}$ coincides with $\mathcal{D}$).
Definition 1.10.
A full triangulated subcategory $\mathcal{C}\subset D^{b}(X)$ is called admissible if the natural inclusion functor $\iota_{\ast}:D^{b}(X)\to D^{b}(X)$ admits both of adjoints $\iota^{\ast}\dashv\iota\dashv\iota^{R}$.
Note that for a triangulated category with the Serre functor, any subcategory that can be a component of semi-orthogonal decomposition of $\mathcal{D}$ is admissible. Conversely, for a admissible subcategory $\mathcal{C}\subset\mathcal{D}$, there are the semi-orthogonal decompositions
$\mathcal{D}=\langle\mathcal{A},\mathcal{C}\rangle=\langle\mathcal{C},\mathcal{B}\rangle.$
Next, we introduce the (full) exceptional collection, which is a most simple form of semi-orthogonal decomposition.
Definition 1.11.
Let $\mathcal{D}$ be a triangulated category and $E$ an object in $\mathcal{D}$. $E$ is called exceptional if the following condition holds:
$\hom^{i}(E,E)=\begin{cases}1&(i=0),\\ 0&(i\neq 0).\end{cases}$
Definition 1.12.
The ordered collection of exceptional objects $(E_{1},\ldots,E_{n})\subset\mathcal{D}$ is called an exceptional collection if for any $i,j\in\{1,\dots,n\}$ with $i<j$ and for any $k\in\mathbb{Z}$,
$\mathop{\mathrm{Hom}}\nolimits^{k}(E_{i},E_{j})=0.$
An exceptional collection $\mathbb{E}=(E_{1},\ldots,E_{n})$ is called
  1. (1)
    full if $\mathbb{E}$ generates the derived category by shifts and extensions;
  2. (2)
    $\operatorname{Ext}$-exceptional if $\mathop{\mathrm{Hom}}\nolimits^{k}(E_{i},E_{j})=0$ for all $i\neq j$, when $k\leq 0$;
  3. (3)
    strong if $\mathop{\mathrm{Hom}}\nolimits^{k}(E_{i},E_{j})=0$ for all $i$ and $j$, when $k\neq 0$.
Example 1.13.
The projective space $\mathbb{P}^{n}$ has a strong full exceptional collection
$D^{b}(X)=\mathbb{E}\coloneqq(\mathscr{O},\mathscr{O}(1),\dots,\mathscr{O}(n)),$
that is called Beilinson Collection.
2. Ulrich Objects: Definition and First Properties
In this Section 2, we generalize the notion of Ulrich bundles to the derived categorical setting by the characterization of Ulrich bundles 1.5. We refer to the basic results about the derived categories to [Huy06]. It will start with the definition of Ulrich objects.
Definition 2.1.
Let $X$ be an algebraic variety of dimension $n$, $\mathscr{O}_{X}(1)$ be a very ample divisor on $X$, and $E$ be an object of the bounded derived category $D^{b}(X)$.
$E$ is said to be an Ulrich object with respect to the polarization $\mathscr{O}_{X}(1)$ if
$H^{i}(E(-j))=0$
for any $i\in\mathbb{Z}$ and $1\leq j\leq n$.
Note that the Ulrich bundle embedded into $D^{b}(X)$ as a complex concentrated in a term is an Ulrich object.
Proposition 2.2.
Let $E\in D^{b}(X)$ be an Ulrich object on a smooth projective variety with an embedding $(X,\mathscr{O}_{X}(1))$ and $\pi:X\to\mathbb{P}^{n}$ a finite linear projection (see 1.4). Then, the following conditions are equivalent:
  1. (1)
    $E\in D^{b}(X)$ is an Ulrich object,
  2. (2)
    $\pi_{\ast}E\in\langle\mathscr{O}_{\mathbb{P}^{n}}\rangle$,
  3. (3)
    each cohomological sheaf $\mathcal{H}^{i}(E)$ is an Ulrich bundle (if not zero).
Proof of 2.2.
Let us prove that (1) implies (2). By the construction of $\pi$, $\pi^{\ast}\mathscr{O}_{\mathbb{P}^{n}}(1)\cong\mathscr{O}_{X}(1)$. Thus, for any $i\in\mathbb{Z}$ and for $j=1,\dots,n$,
$\displaystyle\mathop{\mathrm{Hom}}\nolimits_{D^{b}(\mathbb{P}^{n})}^{i}(\mathscr{O}_{\mathbb{P}^{n}}(j),\pi_{\ast}E)$ $\displaystyle\cong\mathop{\mathrm{Hom}}\nolimits_{D^{b}(X)}^{i}(\pi^{\ast}\mathscr{O}_{\mathbb{P}^{n}}(j),E)$
$\displaystyle\cong\mathop{\mathrm{Hom}}\nolimits_{D^{b}(X)}^{i}(\mathscr{O}_{X}(j),E)$
$\displaystyle\cong H^{i}(E(-j))=0.$
Therefore,
$\pi_{\ast}E\in\langle\mathscr{O}(1),\dots,\mathscr{O}(n)\rangle^{\perp}=\langle\mathscr{O}\rangle,$
that proves what we want.
Next, we will prove that (2) implies (3). As higher direct images of $\mathbf{R}\pi_{\ast}$ vanish since $\pi$ is finite, $\mathbf{R}\pi_{\ast}E$ is the form of
$\dots\to\pi_{\ast}E_{-1}\to\pi_{\ast}E_{0}\to\pi_{\ast}E_{1}\to\cdots\ .$
This complex splits by the assumption (2) and $\langle\mathscr{O}_{\mathbb{P}^{n}}\rangle\cong D^{b}(\operatorname{Spec}\mathbb{C})$, that is,
$\pi_{\ast}E\cong\bigoplus_{i\in\mathbb{Z}}\mathscr{O}_{\mathbb{P}^{n}}^{m_{i}}[-i].$
This shows $\pi_{\ast}\mathcal{H}^{i}(E)\cong\mathscr{O}_{\mathbb{P}^{n}}^{m_{i}}$. Note that here the fact is used that $\mathbf{R}\pi_{\ast}$ commute with $\mathcal{H}^{i}(-)$ (cf. [BvdB03, Corollary 3.3.4.]). Thus, the characterization of Ulrich bundle 1.5 claims that each $\mathcal{H}^{i}(E)$ is Ulrich bundle unless $m_{i}=0$.
Considering the spectral sequence
$E_{2}^{p,q}=H^{p}(\mathcal{H}^{q}(E)(-j))\Longrightarrow H^{p+q}(E(-j)),$
we have that (3) implies (1). ∎
A generalization of the equivalence of (1) and (2) in 2.2 is the following proposition, which is a well-known statement for the case of Ulrich bundles.
Proposition 2.3.
Let $\pi:X\to Y$ be a finite morphism between the same dimensional smooth projective varieties, and $\mathscr{O}_{Y}(1)$ be a very ample line bundle on $Y$. Assume that $\pi^{\ast}\mathscr{O}_{Y}(1)$ is also very ample on $X$. Then, $E$ is Ulrich object with respect to $\pi^{\ast}\mathscr{O}_{Y}(1)$ if and only if $\pi_{\ast}E$ is Ulrich bundle with respect to $\mathscr{O}_{Y}(1)$.
Proof of 2.3.
Considering the Leray’s spectral sequence
$E_{2}^{p,q}=H^{q}(Y,R^{p}\pi_{\ast}E(-j))\Rightarrow H^{p+q}(X,E(-j))$
and higher direct image vanishes because $\pi$ is finite, we have the claim. ∎
Proposition 2.4.
Let $Y\in|H|$ be a general hyperplane section of $X$, $E$ an Ulrich object on $X$, and $\mathscr{O}_{Y}(1)\coloneqq\mathscr{O}_{X}(1)|_{Y}$. Then, $E|_{Y}$ is also an Ulrich object for $(Y,\mathscr{O}_{Y}(1))$.
Proof of 2.4.
Considering the triangle
$E(-j-1)\to E(-j)\to E|_{Y}(-j)\to E(-j-1)[1]$
and setting $j=1,\dots,n-1$, $H^{i}(E|_{Y}(-j))=0$ hold for any $i\in\mathbb{Z}$ and $j=1,\dots,n-1$. ∎
Lemma 2.5.
Let $E,F,G$ be objects of $D^{b}(X)$ such that there exists the distinguished triangle
$E\to F\to G\to E[1].$
If two of $E,F,G$ are Ulrich objects, then the other is.
Proof of 2.5.
For each $j=1,\dots,n$,
$E(-j)\to F(-j)\to G(-j)\to E(-j)[1]$
is a distinguished triangle. Thus, the claim can be concluded by taking the long exact sequence. ∎
Proposition 2.6.
Let $E$ (resp. $F$) be an Ulrich object on $(X,\mathscr{O}_{X}(1))$ (resp. $(Y,\mathscr{O}_{Y}(1))$). Then, $E\boxtimes F(\dim X)$ and $E(\dim Y)\boxtimes F$ are Ulrich objects on $(X\times Y,\mathscr{O}_{X}(1)\boxtimes\mathscr{O}_{Y}(1))$.
Proof of 2.6.
Let $n=\dim X$. Let us show the case of $E\boxtimes F(\dim X)$. The claim easily follows from the Künneth formula
$H^{i}(X\times Y,E\boxtimes F(n)\otimes\mathscr{O}_{X}(-j)\boxtimes\mathscr{O}_{Y}(-j))=\bigoplus_{p+q=i}H^{p}(X,E(-j))\otimes H^{q}(Y,F(n-t)).$
3. Ulrich Objects: Examples
3.1. Yoneda Type Ulrich Objects
The next subject of consideration shifts to the Yoneda extension. We refer [Oor64] for the basic notions and results of the Yoneda extension.
Definition 3.1.
Let $\mathcal{F}$ and $\mathcal{G}$ be sheaves on $X$. Assume that there exists a non-trivial element $\eta\in\operatorname{Ext}^{k}(\mathcal{F},\mathcal{G})$ for some $k\geq 1$. Then, there is the following exact sequence corresponding to $\eta$:
$0\to\mathcal{G}\to E_{k-1}\to\cdots\to E_{0}\to\mathcal{F}\to 0$
In this paper, we refer to the object
$E_{\eta}:(\cdots 0\to E_{k-1}\to\cdots\to E_{0}\to 0\cdots)\in D^{b}(X)$
as the Yoneda extension of degree $k$ corresponds to $\eta\in\operatorname{Ext}^{k}(\mathcal{F},\mathcal{G})$ (here, the degree of $E_{j}$ as the complex is the same as $j$).
Note that two Yoneda extensions
$0\to\mathcal{G}\to E_{k-1}\to\cdots\to E_{0}\to\mathcal{F}\to 0,$
and
$0\to\mathcal{G}\to F_{k-1}\to\cdots\to F_{0}\to\mathcal{F}\to 0$
are said to be equivalent if there exists another Yoneda extension
$0\to\mathcal{G}\to G_{k-1}\to\cdots\to G_{0}\to\mathcal{F}\to 0$
such that there is the following commutative diagram:
[…]
The Yoneda extension $E$ is isomorphic to $\mathcal{G}[-k+1]\oplus\mathcal{F}$ if and only if $E$ is corresponding to $0\in\operatorname{Ext}^{k}(\mathcal{F},\mathcal{G})$. In particular, it is observed that the Yoneda extension gives a non-trivial example of Ulrich object:
Example 3.2.
Let $\mathcal{F}$ and $\mathcal{G}$ be Ulrich bundles. Assume that there is a non-trivial element $\eta\in\operatorname{Ext}^{m}(\mathcal{F},\mathcal{G})$ for $m\geq 1$. Then, Yoneda extension defined in 3.1
$E_{\eta}:\cdots\to 0\to E_{-m+1}\to\cdots\to E_{0}\to 0\to\cdots$
correspond to $\eta$ is an Ulrich object.
Indeed, because
$\mathcal{H}^{i}(E)=\begin{cases}\mathcal{F}&\text{if }i=0,\\ \mathcal{G}&\text{if }i=-m+1,\\ 0&\text{otherwise},\end{cases}$
it is shown from 2.2.
Definition 3.3.
We call the Ulrich object constructed in the way of 3.2 (up to shifts) as Yoneda type Ulrich object.
3.2. Explicit Examples
First, directly from 2.2, the object $E\in D^{b}(\mathbb{P}^{n})$ is Ulrich with respect to $\mathscr{O}(1)$ if and only if $E\in\langle\mathscr{O}\rangle$.
When the dimension of the variety is 11 11, Ulrich objects are just a direct sum of Ulrich bundles from 2.2. Indeed, in this case, any object $E$ is formal, that is, $E$ can be written as
$E\cong\bigoplus_{i\in\mathbb{Z}}\mathcal{H}^{i}(\mathcal{E})[-i]$
(see [Huy06, Corollary 3.15]). Thus, any indecomposable Ulrich object on $(C,\mathscr{O}(1))$ is only an Ulrich bundle.
3.2.1. Ulrich Objects on Quadric Hypersurfaces
The Ulrich bundles on quadric hypersurfaces $Q_{n}$ are bundles known as Spinor bundles. The original reference of spinor bundles on quadrics is [Ott88]. The derived categorical treatment of quadrics are in [Kap88, Section 4.].
Theorem 3.4 ([Kap88]).
Let $Q_{n}\subset\mathbb{P}^{n+1}$ be a quadric hypersurface. Then, there is a full exceptional collection
$D^{b}(Q_{n})=\begin{cases}\langle\mathscr{O},S,\mathscr{O}(1),\dots,\mathscr{O}(n-1)\rangle&\text{if }n:\text{odd,}\\ \langle\mathscr{O},S^{+},S^{-},\mathscr{O}(1),\dots,\mathscr{O}(n-1)\rangle&\text{if }n:\text{even,}\end{cases}$
where $S$, $S^{+}$, and $S^{-}$ are the spinor bundles.
The Kapranov collection above claims that $E\in D^{b}(Q_{n})$ is Ulrich object if and only if $E\in\langle S\rangle$ when $n$ is odd and $E\in\langle S^{+},S^{-}\rangle$ when $n$ is even.
More precisely, when $n$ is odd Ulrich object $E$ is of the form of
$E\cong\bigoplus_{i\in\mathbb{Z}}S^{m_{i}}[-i].$
Noting that $S^{+}$ and $S^{-}$ are orthogonal, when $n$ is even Ulrich object $E$ is of the form of
$E\cong\bigoplus_{i\in\mathbb{Z}}(S^{-})^{m_{i}}[-i]\oplus\bigoplus_{j\in\mathbb{Z}}(S^{+})^{m_{j}}[-j].$
3.2.2. Ulrich Objects on Surfaces
In the case of curve $C$, any object in $D^{b}(C)$ splits into the direct sum of shifts of sheaves, but in the case of surface, the following proposition is known by Uehara and Ishii:
Proposition 3.5 ([IU05, Proposition 3.2.]).
Let $S$ be a surface. Giving a object in $D^{b}(C)$ is equivalent to giving finitely many sheaves $\mathcal{G}^{i}$ and element
$e^{i}\in\operatorname{Ext}^{2}(\mathcal{G}^{i},\mathcal{G}^{i-1})$
for each $i$.
Thus, giving an Ulrich object on a surface $S$ is equivalent to giving finitely many Ulrich sheaves $\mathcal{G}^{i}$ and their degree 22 22 extensions.
Moreover, the Ulrichness of the object imposes a strong restriction on its Chern character. To see this, for example, let $X$ be a surface of Picard rank 11 11. Let us write $\operatorname{Pic}(X)=\mathbb{Z}[H]$, $H^{2}=d$, and $K_{X}=i_{X}H$.
Lemma 3.6.
Let $E$ be an Ulrich object of rank $r$ on $X$. Then,
$\displaystyle\operatorname{ch}([E])=r+\frac{r}{2}(i_{X}+3)H+\bigg{(}-r\chi(\mathscr{O}_{X})+\frac{rd}{4}(i_{X}^{2}+3i_{X}+4)\bigg{)}.$ (3.1)
Proof of 3.6.
As $\rho(X)=1$, we can write $\operatorname{ch}_{1}E=e_{1}H$ and $\operatorname{ch}_{2}E=e_{2}H^{2}=e_{2}d[\operatorname{pt}]$ with $e_{1},e_{2}\in\mathbb{Z}$. Then,
$\chi(E(-k))=r+(e_{1}-rkH)+\bigg{(}e_{2}d-e_{1}kH+\frac{k^{2}}{2}rd\bigg{)}$
and the Hirzebruch-Riemann-Roch Theorem induce
$\displaystyle r\chi(\mathscr{O}_{X})+d\bigg{(}-\bigg{(}\frac{i_{X}}{2}+1\bigg{)}e_{1}+e_{2}+\frac{r}{2}\big{(}i_{X}+1\big{)}\bigg{)}$ $\displaystyle=0,$
$\displaystyle r\chi(\mathscr{O}_{X})+d\bigg{(}-\bigg{(}\frac{i_{X}}{2}+2\bigg{)}e_{1}+e_{2}+r\big{(}i_{X}+2\big{)}\bigg{)}$ $\displaystyle=0.$
Thus, we have the result. ∎
4. Application to the Eisenbud-Schreyer’s Question
This Section 4 states the application to the Eisenbud-Schreyer conjecture to show the evidence of the merit for considering the Ulrich object. As previously stated, for the sake of convenience, the claim of the Eisenbud-Schreyer’s conjecture A is reiterated here: any smooth projective variety carries at least an Ulrich bundle.
4.1. Generators of the Derived Category
For the notions of the generator of the derived category, see [BvdB03] or [Sta18, Tag 09SI] for more details.
Before proceeding to the definition of generators, a notation is introduced:
$\displaystyle\langle\Omega\rangle^{\perp}$$\displaystyle\coloneqq\{E\in\mathcal{D}\mid\mathop{\mathbf{R}\mathrm{Hom}}\nolimits(X,E)=0,{}^{\forall}X\in\langle\Omega\rangle\}$
=$\displaystyle=\{E\in\mathcal{D}\mid\mathop{\mathbf{R}\mathrm{Hom}}\nolimits(X,E)=0,{}^{\forall}X\in\Omega\}.$
Definition 4.1.
Let $\mathcal{D}$ be a triangulated category and $G$ be an object of $\mathcal{D}$. Then, $G$ is
  1. (1)
    classical generator if $\langle G\rangle=\mathcal{D}$, and
  2. (2)
    weak generator (or simply, generator) if $\langle G\rangle^{\perp}=0$.
As a remark, we mention the notion of strong generator. $G$ is said to be a strong generator if the entire derived category can be generated in a finite number of steps by operations involving direct summands and extensions, starting with $G$. (See [BvdB03] or [Sta18, Tag 09SI] for the precise definition.) By definition, a strong generator is a classical generator. For the other direction of this implication, the following is known:
Lemma 4.2 ([Sta18, Tag 0FXA]).
Let $\mathcal{D}$ be a triangulated category that has a strong generator. Let $E$ be an object of $\mathcal{D}$. If $E$ is a classical generator of $\mathcal{D}$, then $E$ is a strong generator.
Thus, we need not distinguish between classical and strong in our cases.
Theorem 4.3 (see [BvdB03, Theorem 3.1.4.]).
Let $X$ be a smooth projective variety. Then, $D^{b}(X)$ admits a strong generator.
The following Theorem says that polarizing a very ample line bundle on $X$ fixes a classical generator of $D^{b}(X)$ in a sense.
Theorem 4.4 ([Orl09, Theorem 4.]).
Let $X$ be a smooth projective variety and $\mathscr{O}_{X}(1)$ be a very ample line bundle on $X$. Then, $G=\mathscr{O}_{X}\oplus\mathscr{O}_{X}(1)\oplus\dots\oplus\mathscr{O}_{X}(n)$ is a classical generator on $X$.
Proposition 4.5 ([Sta18, Tag 09SR], [Nee92]).
Let $\mathcal{D}$ be a triangulated category with direct sums and $G$ be a compact object of $\mathcal{D}$. Then, the following are equivalent:
  1. (1)
    $G$ is a classical generator for $\mathcal{D}^{\mathrm{cpt}}$ and $\mathcal{D}$ is compactly generated, and
  2. (2)
    $G$ is a generator for $\mathcal{D}$.
4.2. Existence for Curves
In this Section 4.2, we will show D.
Lemma 4.6.
Let $\mathcal{D}$ be a triangulated category with arbitrary direct sums and $G$ a weak generator of $\mathcal{D}^{\mathrm{cpt}}$. Let $f:A\to B$ be a morphism in $D$ such that $C\coloneqq\operatorname{Cone}f\cong\bigoplus_{i}E_{i}$, where $E_{i}\in\mathcal{D}^{\mathrm{cpt}}(X)$. Then,
$f\circ-:\mathop{\mathbf{R}\mathrm{Hom}}\nolimits(G,A)\stackrel{{\scriptstyle\cong}}{{\to}}\mathop{\mathbf{R}\mathrm{Hom}}\nolimits(G,B)$
if and only if $f$ is an isomorphism.
Proof of 4.6.
By applying $\mathop{\mathbf{R}\mathrm{Hom}}\nolimits(G,-)$ to the triangle defined by $f$, the following triangle arises:
$\mathop{\mathbf{R}\mathrm{Hom}}\nolimits(G,A)\to\mathop{\mathbf{R}\mathrm{Hom}}\nolimits(G,B)\to\mathop{\mathbf{R}\mathrm{Hom}}\nolimits(G,C)\cong\bigoplus_{i}\mathop{\mathbf{R}\mathrm{Hom}}\nolimits(G,E_{i})\stackrel{{\scriptstyle[1]}}{{\to}}.$
Thus, if $f\circ-:\mathop{\mathbf{R}\mathrm{Hom}}\nolimits(G,A)\stackrel{{\scriptstyle\cong}}{{\to}}\mathop{\mathbf{R}\mathrm{Hom}}\nolimits(G,B)$, we have $\mathop{\mathbf{R}\mathrm{Hom}}\nolimits(G,E_{i})=0$ for all $i$. As $G$ weakly generates $\mathcal{D}^{\mathrm{cpt}}$, we have the result. ∎
In the proof of following 4.9, the homotopy colimit is used to approximate an object $X$ in a triangulated category with arbitrary direct sums by a set of objects in $\mathcal{D}^{\mathrm{cpt}}$ which is not necessarily finite.
Definition 4.7.
Let $\mathcal{D}$ be a triangulated category and $(X_{n},j_{n}:X_{n-1}\to X_{n})_{n\in\mathbb{Z}_{\geq 0}}$ be a system of $\mathcal{D}$. A homotopy colimit of the system $(X_{n})$ is defined as the objects fitting into the following triangle:
$\bigoplus_{n}X_{n}\stackrel{{\scriptstyle 1-j}}{{\longrightarrow}}\bigoplus_{n}X_{n}\to X\stackrel{{\scriptstyle[1]}}{{\to}}\Big{(}\bigoplus_{n}X_{n}\Big{)}[1],$
where
$1-j\coloneqq\begin{pmatrix}1&0&0&\cdots&&\\ -j_{1}&1&0&&&\\ 0&-j_{2}&1&&&\\ &\vdots&&\ddots&&\end{pmatrix}.$
Write this homotopy colimit $X$ by $\operatorname{hocolim}X_{n}$.
Lemma 4.8 ([Sta18, Tag 09SN]).
Let $\mathcal{D}$ be a triangulated category with direct sums and $\{E_{i}\}_{i\in I}\subset\operatorname{ob}(\mathcal{D}^{\mathrm{cpt}})$ objects such that generates $\mathcal{D}$. Then, for any $K\in\mathcal{D}$ there exists a sequence
$X_{1}\stackrel{{\scriptstyle j_{1}}}{{\to}}X_{2}\stackrel{{\scriptstyle j_{2}}}{{\to}}X_{3}\stackrel{{\scriptstyle j_{3}}}{{\to}}\cdots$
such that $X_{1}$ and each $\operatorname{Cone}(j_{i})$ are the direct sum of shifts of the $E_{i}$ and
$K=\operatorname{hocolim}X_{n}.$
Proposition 4.9.
Let $C$ be a curve and $G\in D^{b}(C)$. If $G$ weakly generates $D^{b}(C)$, $G$ classically generates $D^{b}(C)$.
Proof of 4.9.
Let $K\in D(\operatorname{\mathbf{Qcoh}}C)$ and assume that $\mathop{\mathbf{R}\mathrm{Hom}}\nolimits(G,K)=0$. By the 4.8, there exists a sequence
$K_{1}\stackrel{{\scriptstyle j_{1}}}{{\to}}K_{2}\stackrel{{\scriptstyle j_{2}}}{{\to}}K_{3}\stackrel{{\scriptstyle j_{3}}}{{\to}}\cdots$
such that $K=\operatorname{hocolim}K_{n}$ and $\operatorname{Cone}(j_{i})=\bigoplus_{k}\mathcal{E}_{k}^{i}[-k]$,where $\mathcal{E}_{j}^{i}\in\operatorname{\mathbf{Coh}}(C)$. Here, for the convenience, write $\operatorname{Cone}(j_{0})=K_{1}$.
As any object in $D^{b}(C)$ splits, coherent sheaves $\mathcal{F}_{j}$ can rewrite $K_{i}$ as
$K_{i}\cong\bigoplus_{k}\mathcal{F}_{k}[-k].$
Thus, $K=\operatorname{Cone}(1-j)$ splits into the (infinite) direct sum of the shifts of the coherent sheaves. 4.6 claims that the $1-j$ is an isomorphism and therefore, $K\cong 0$. By 4.5, $G$ classically generates $D^{b}(C)$. ∎
Theorem 4.10.
Let $\Phi_{|\mathscr{O}_{C}(1)|}:C\hookrightarrow\mathbb{P}^{N}$ be a smooth projective curve with an embedding. Then, an Ulrich object exists on $(C,\mathscr{O}_{C}(1))$.
Proof of 4.10.
Note that $\mathscr{O}(1)\in D^{b}(C)$ cannot classically generate the whole derived cateogy $D^{b}(C)$. Indeed, as $\operatorname{rk}K_{0}(\langle\mathscr{O}(1)\rangle)=1$ by the stability of $\mathscr{O}(1)$, comparing the rank of the K-group of $\langle\mathscr{O}(1)\rangle$ and $D^{b}(C)$, we have
$1\displaystyle 2=\operatorname{rk}K_{0}(D^{b}(C))>\operatorname{rk}K_{0}(\langle\mathscr{O}_{C}(1)\rangle)=$ 1 .
Thus, $\mathscr{O}(1)$ cannot classically generate the whole derived category. Finally, 4.9 shows the claim. ∎
5. Remarks on Bridgeland Stability of Ulrich Objects
The last section is devoted to some remarks on the relationship between the Ulrich objects and stabilities in the derived category, namely the Bridgeland stabilities.
As stated in Section 1.1, the Ulrich bundles behave well for the stabilities of sheaves. Thus, it predicts that the Ulrich objects behave well for Bridgeland stabilities. However, in general, investigating the Bridgeland stabilities is not easy.
5.1. Preliminaries on Bridgeland Stability Conditions
Definition 5.1.
Let $\mathcal{D}$ be a triangulated category, $\mathcal{A}$ an abelian category and $K_{\mathrm{num}}(\mathcal{A})$ and $K_{\mathrm{num}}(\mathcal{A})$ be these numerical Grothendieck group respectively.
  • A pre-stability function on $\mathcal{A}$ is a group homomorphism $Z:K_{\mathrm{num}}(\mathcal{A})\to\mathbb{C}$ such that for all $A\in\mathcal{A}\setminus 0$, $Z(A)\in\mathbb{H}^{+}(=\mathbb{H}\cup\mathbb{R}_{<0})$.
  • A stability function $Z$ on $\mathcal{A}$ is a pre-stability function on $\mathcal{A}$ which satisfies the Harder-Narasimhan property.
  • A Bridgeland stability condition $\sigma$ on $\mathcal{D}$ is a pair $(\mathcal{A},Z)$, where $\mathcal{A}\subset\mathcal{D}$ is an abelian category and $Z$ is a stability function on $\mathcal{A}$.
Denote the set of Bridgeland stability conditions on $\mathcal{D}$ by $\operatorname{Stab}(\mathcal{D})$. Also, for a smooth projective variety $X$, write $\operatorname{Stab}(X)\coloneqq\operatorname{Stab}(D^{b}(X))$ for simplicity. Note that in [Bri07] Bridgeland have shown that $\operatorname{Stab}(X)$ become a complex manifold with a suitable topology.
Proposition 5.2.
Let $X$ be a surface, $D,H$ be an $\mathbb{R}$-divisors such that $H$ is ample. Define an abelian category by
$\mathcal{A}_{D,H}=\{F\in D^{b}(X)\ \mid\ H^{i}(F)=0\ \text{for}\ i\neq-1,0,\ H^{-1}(F)\in\mathcal{F}_{D,H},H^{0}(F)\in\mathcal{T}_{D,H}\},$
where the pair $(\mathcal{F}_{D,H},\mathcal{T}_{D,H})$ is the torsion pair defined by
$\mathcal{F}_{D,H}=\{E\in\operatorname{\mathbf{Coh}}(X)\ \mid\ E:\text{torsion free and}\ \forall 0\neq F\subset E:\mu_{H}(F)\leq D.H\},\text{and}$
$\mathcal{T}_{D,H}=\{E\in\operatorname{\mathbf{Coh}}(X)\ \mid\ \forall E\twoheadrightarrow F\neq 0\text{ torsion free}:\ \mu_{H}(F)>D.H\},$
where $\mu_{H}(E)=\frac{c_{1}(E).H}{\operatorname{rk}(E)H^{2}}$ is the slope function. For $F\in\mathcal{A}_{D,H}$, set
$\displaystyle Z_{D,H}(F)=-\int\exp(-(D+iH))\operatorname{ch}(F).$
Then, there exists the injection:
$\displaystyle N^{1}(X)\times\operatorname{Amp}(X)$ $\displaystyle\longrightarrow\operatorname{Stab}(X)$
$\displaystyle(D,H)\hskip 28.45274pt$ $\displaystyle\longmapsto\sigma_{D,H}=(Z_{D,H},\mathcal{A}_{D,H}).$
Definition 5.3.
Let $\mathcal{D}$ be a triangulated category. A stability condition $\sigma=(Z,\mathcal{A})$ is called geometric if skyscraper sheaf $k(x)$ is $\sigma$-semistable of same phase for any $x\in X$.
Proposition 5.4.
Let $X$ be a smooth projective surface over $\mathbb{C}$. Then, the divisorial stability condition $\sigma_{D,H}$ is geometric such that the phase of the skyscraper is 11 11. Conversely, if $\sigma\in\operatorname{Stab}(X)$ satisfies that $k(x)$ is $\sigma$-stable of phase 11 11 for any $x\in X$, then there exists a pair $(D,H)\in N^{1}(X)_{\mathbb{R}}\times\operatorname{Amp}(X)_{\mathbb{R}}$ such that $\sigma=\sigma_{D,H}$.
5.2. First Examples
Example 5.5.
Let $C$ be a smooth projective curve of genus $g>0$ and $\sigma$ be a Bridgeland stability condition on $D^{b}(C)$. An Ulrich object $E$ is trivially semistable from 1.6.
5.2.1. The Projective Plane
From 5.2, the divisorial stability condition in $\operatorname{Stab_{\mathrm{div}}}(\mathbb{P}^{2})$ is determined by the pair in the upper half-plane $(s,t)\in\mathbb{H}$. Denote $\sigma_{sH,tH}$ by $\sigma_{s,t}$.
Theorem 5.6 ([AM16, Theorem 5.4.]).
Let $S$ be a surface that does not have any negative self-intersection curve, and $\sigma=(Z,\mathcal{A})\in\operatorname{Stab_{\mathrm{div}}}(X)$ such that $\mathscr{O}_{S}\in\mathcal{A}$. Then, $\mathscr{O}_{S}$ is stable with respect to $\sigma$.
In particular, the structure sheaf $\mathscr{O}_{\mathbb{P}^{2}}$ of the projective plane $\mathbb{P}^{2}$ is Bridgeland stable for any $\sigma_{sH,tH}$ $(s<0)$.
The dual version of the 5.6 is the following:
Proposition 5.7 ([AM16, Proposition 6.3.]).
Let $S$ be a surface that does not have any negative self-intersection curve, and $\sigma=(Z,\mathcal{A})\in\operatorname{Stab_{\mathrm{div}}}(X)$ such that $\mathscr{O}_{S}[1]\in\mathcal{A}$. Then, $\mathscr{O}_{S}[1]$ is stable with respect to $\sigma$.
Thus, the shifted structure sheaf $\mathscr{O}_{\mathbb{P}^{2}}[1]$ is Bridgeland stable for any $\sigma_{sH,tH}$ $(s>0)$.
As a corollary of the results in [AM16], we have the following statement:
Proposition 5.8.
Let $E$ be an Ulrich object on $\mathbb{P}^{2}$. Assume that there exists an integer $k\in\mathbb{Z}$ and a divisorial stability condition $\sigma_{s,t}=(Z_{s,t},\mathcal{A}_{s,t})$ on $D^{b}(\mathbb{P}^{2})$ such that $E[k]\in\mathcal{A}_{s,t}$. Then, $E$ is $\sigma_{s,t}$-semistable.
Proof of 5.8.
Arcara-Miles says $\mathscr{O}_{\mathbb{P}^{2}}$ is $\sigma_{s,t}$-stable. An Ulrich object on $\mathbb{P}^{2}$ is the form of
$E\cong\bigoplus_{i\in\mathbb{Z}}\mathscr{O}_{\mathbb{P}^{2}}^{n_{i}}[-i].$
by 2.2. By the definition of $\mathcal{A}_{s,t}$, only one $n_{i}\neq 0$ if $E[k]\in\mathcal{A}_{s,t}$. ∎
5.2.2. Quadrics
The Spinor bundles on the quadric surface $Q^{2}\cong\mathbb{P}^{1}\times\mathbb{P}^{1}$ are $\mathscr{O}(1,0)$ and $\mathscr{O}(0,1)$ via the identification. Thus, Spinor bundles on $Q^{2}$ is $\sigma$-stable for $\sigma\in\operatorname{Stab}_{\mathrm{div}}(Q^{2})$ up to shift from [AM16, Theorem 5.4. and Proposition 6.3.].
Next, Let $Q^{3}$ be the quadric threefold. The Spinor bundle $S$ is an object in the Kuznetsov component $\operatorname{Ku}(Q^{3})$. [Yan24, Lemma A.1.] claims that the Spinor bundle is $\sigma$-stable for any
$\sigma\in\mathcal{K}\coloneqq\Big{\{}\sigma\left(\alpha,-\frac{1}{2}\right)\cdot\widetilde{\operatorname{GL}}_{2}^{+}(\mathbb{R})\Big{\}}\subset\operatorname{Stab}(\operatorname{Ku}(Q^{3})).$
See [Yan24, Section 4.] for more precise definitions.
5.2.3. Bridgeland Stability of Ulrich Objects on Surfaces
Proposition 5.9.
Let $X$ be a surface and $E\in D^{b}(X)$ be an Ulrich object with respect to an embedding $\mathscr{O}_{X}(1)$.
Assume that $E$ is contained in a tilting heart $\mathcal{A}_{D,H}$ for some $(D,H)\in N^{1}(X)_{\mathbb{R}}\times\operatorname{Amp}(X)_{\mathbb{R}}$. Then, $E$ is a shift of an Ulrich sheaf.
Proof of 5.9.
Assume $\mathcal{H}^{i}(E)\neq 0$ for $i=-1,0$. Note that $\mathcal{H}^{i}(E)$ are both Ulrich bundles by 2.2 and semistable by 1.6. As the slopes of $\mathcal{H}^{i}(E)$ coincide, it contradict to the construction of $\mathcal{A}_{D,H}=\langle\mathcal{F}_{D,H}[1],\mathcal{T}_{D,H}\rangle$. ∎
Question 5.10.
Let $X$ be a smooth surface over $\mathbb{C}$. Is there any geometric stability condition $\sigma=(Z,\mathcal{A})\in\operatorname{Stab}(X)$ and an Ulrich object $E$ such that $E\in\mathcal{A}$ that is not $\sigma$-semistable?